A regression was run to determine if there is a relationship between hours of TV watched per day ( x x ) and number of situps a person can do ( y y ). The results of the regression were: y=ax+b

a=-1.154

b=23.666 r2=0.473344

r=-0.688

Use this to predict the number of situps a person who watches 5 hour(s) of TV can do, and please round your answer to a whole number.
Answer #1

x: hours of TV watched per day

y: number of situps a person can do

y=ax+b

where a=-1.154 b=23.666

The regression equation is

y = -1.154 + 23.666 (x)

predict the number of situps a person who watches 5 hour(s) of TV can do

i.e. x =5 hours

y= -1.154 + 23.666(5)

= -1.154 + 118.33

= 117.18 ~~ 117

117 situps a person can do when he watches 5 hour(s) of TV